# SCoT: Straight Consistent Trajectory for Pre-Trained Diffusion Model Distillations

### v3

@@ 4.2 Distillation Results
CIFAR-10. On CIFAR-10, SCoT achieves an FID of 2.30 with only 2 NFEs, outperforming baselines such as CD (FID 2.93) and CT (FID 5.83) while requiring fewer function evaluations. CTM attains a lower FID of 1.87 with just 1 NFE, but its results incorporate a GAN loss, which significantly improves sample fidelity by providing a strong adversarial signal during training. In contrast, we do not adopt the GAN loss due to its instability under our training configuration and resource constraints. Despite this, SCoT delivers competitive performance without adversarial training, demonstrating a favorable balance between generation quality and computational efficiency. Notably, in the unconditional setting, SCoT (FID 2.30) also surpasses CM (FID 3.55), further highlighting its effectiveness in fast, high-quality image generation.
ImageNet. On ImageNet, SCoT achieves an FID of 2.60 with just 2 NFEs, outperforming baselines such as CD (FID 6.20) and PD (FID 15.39), and closely approaching the performance of EDM (FID 2.44). While CTM achieves the best FID of 1.92, this result benefits from the integration of a GAN loss, which enhances visual fidelity by introducing adversarial supervision during training. In contrast, our method does not incorporate GAN loss due to its instability under our training regime, yet still attains competitive sample quality. SCoT also demonstrates strong diversity, with a recall of 0.61 and an Inception Score of 68.2, highlighting its effectiveness in balancing quality, diversity, and computational cost.
@@ 4.5 Velocity Guarantee
The incorporation of the velocity guarantee loss, as shown in Table˜5, further improves the generative trajectory by encouraging temporal smoothness. This constraint ensures that the trajectories are straightened, making the sampling process more stable and efficient. The benefits are evident in the substantial reduction of FID, dropping from 14.7 to 4.8 in Step=1 (NFE=2) and from 15.2 to 3.6 in Step=2. The IS improves accordingly, reaching 67.6 and 68.2 in Step=1 and Step=2, respectively. The model also achieves stronger coverage and accuracy, with Precision improving from 0.56 to 0.71 and Recall from 0.54 to 0.58 in Step=1. Similar trends are observed in Step=2. These confirm that velocity regularization helps refine the generative path, enabling the model to produce high-quality samples with better alignment to the underlying data manifold, especially under reduced NFE.
@@
Table 4: Comparison of training results for NFE=1 and NFE=2 for loss from Equation˜3. All models are trained on ImageNet and tested on 6k samples. Bold values indicate the best performance per metric.
@@
Table 5: Training Equation˜4 Results for NFE=1 and NFE=2. For metrics with $\downarrow$, lower values are better; for metrics with $\uparrow$, higher values are better. Bold values represent the best performance in each metric.
@@ 4.6 Loss Weighting
In this section, we explore the impact of loss weighting parameters in our training objective defined by Equation˜4, focusing specifically on $\lambda_{\text{con}}$ for the consistency loss and $\lambda_{\text{vel}}$ for the velocity loss.
Weighting on $\lambda_{\text{con}}$ in Equation˜4 We implemented an adaptive weighting strategy for $\lambda_{\text{con}}$ to effectively balance the contributions between consistency and denoising losses during training. The primary motivation behind this adaptive mechanism is to dynamically adjust the weighting based on the relative importance and magnitude of each loss component. Specifically, this adjustment is made through gradient magnitude comparisons, ensuring that no single loss dominates excessively, thereby stabilizing training and improving convergence.
Weighting on $\lambda_{\text{vel}}$ in Equation˜4  To address the instability associated with the velocity loss defined in Equation˜2, we explored several loss weighting strategies. The primary source of instability stems from the computation of second-order derivatives using automatic differentiation tools such as PyTorch’s torch.autograd.grad(). These second-order computations require repeated differentiation, which can amplify small numerical inaccuracies through successive applications of the chain rule. As a result, accumulated numerical errors may lead to unstable gradient magnitudes, manifesting as gradient explosion or vanishing gradients, and ultimately impairing training stability. Motivated by these challenges, we investigated alternative weighting approaches to mitigate such instabilities.
@@ D.1 Time Efficiency
Dataset Equation˜2 Equation˜3

### v4

@@ 4.2 Distillation Results
CIFAR-10.
On CIFAR-10, SCoT achieves an FID of 2.30 with only 2 NFEs, outperforming baselines such as CD (FID 2.93) and CT (FID 5.83) while requiring fewer function evaluations. CTM attains a lower FID of 1.87 with just 1 NFE, but its results incorporate a GAN loss, which significantly improves sample fidelity by providing a strong adversarial signal during training. In contrast, we do not adopt the GAN loss due to its instability under our training configuration and resource constraints. Despite this, SCoT delivers competitive performance without adversarial training, demonstrating a favorable balance between generation quality and computational efficiency. Notably, in the unconditional setting, SCoT (FID 2.30) also surpasses CM (FID 3.55), further highlighting its effectiveness in fast, high-quality image generation.
ImageNet.
On ImageNet, SCoT achieves an FID of 2.60 with just 2 NFEs, outperforming baselines such as CD (FID 6.20) and PD (FID 15.39), and closely approaching the performance of EDM (FID 2.44). While CTM achieves the best FID of 1.92, this result benefits from the integration of a GAN loss, which enhances visual fidelity by introducing adversarial supervision during training. In contrast, our method does not incorporate GAN loss due to its instability under our training regime, yet still attains competitive sample quality. SCoT also demonstrates strong diversity, with a recall of 0.61 and an Inception Score of 68.2, highlighting its effectiveness in balancing quality, diversity, and computational cost.
@@ 4.5 Velocity Guarantee
The incorporation of the velocity guarantee loss, as shown in Table˜5, further improves the generative trajectory by encouraging temporal smoothness. This constraint ensures that the trajectories are straightened, making the sampling process more stable and efficient. The benefits are evident in the substantial reduction of FID, dropping from 14.7 to 4.8 in Step=1 (NFE=2) and from 15.2 to 3.6 in Step=2 (NFE = 2). The IS improves accordingly, reaching 67.6 and 68.2 in Step=1 (NFE = 1) and Step=2 (NFE = 2), respectively. The model also achieves stronger coverage and accuracy, with Precision improving from 0.56 to 0.71 and Recall from 0.54 to 0.58 in Step=1 (NFE = 1). Similar trends are observed in Step=2 (NFE = 2). These confirm that velocity regularization helps refine the generative path, enabling the model to produce high-quality samples with better alignment to the underlying data manifold, especially under reduced NFE.
@@
Table 4: Comparison of training results for NFE=1 and NFE=2 for loss from Equation˜4. All models are trained on ImageNet and tested on 6k samples. Bold values indicate the best performance per metric.
@@
Table 5: Training Equation˜5 Results for NFE=1 and NFE=2. For metrics with $\downarrow$, lower values are better; for metrics with $\uparrow$, higher values are better. Bold values represent the best performance in each metric.
@@ 4.6 Loss Weighting
In this section, we explore the impact of loss weighting parameters in our training objective defined by Equation˜5, focusing specifically on $\lambda_{\text{con}}$ for the consistency loss and $\lambda_{\text{vel}}$ for the velocity loss.
Weighting on $\lambda_{\text{con}}$ in Equation˜5 We implemented an adaptive weighting strategy for $\lambda_{\text{con}}$ to effectively balance the contributions between consistency and denoising losses during training. The primary motivation behind this adaptive mechanism is to dynamically adjust the weighting based on the relative importance and magnitude of each loss component. Specifically, this adjustment is made through gradient magnitude comparisons, ensuring that no single loss dominates excessively, thereby stabilizing training and improving convergence.
Weighting on $\lambda_{\text{vel}}$ in Equation˜5  To address the instability associated with the velocity loss defined in Equation˜2, we explored several loss weighting strategies. The primary source of instability stems from the computation of second-order derivatives using automatic differentiation tools such as PyTorch’s torch.autograd.grad(). These second-order computations require repeated differentiation, which can amplify small numerical inaccuracies through successive applications of the chain rule. As a result, accumulated numerical errors may lead to unstable gradient magnitudes, manifesting as gradient explosion or vanishing gradients, and ultimately impairing training stability. Motivated by these challenges, we investigated alternative weighting approaches to mitigate such instabilities.
@@ 5 Conclusion
6 Acknowledgments and Disclosure of Funding
This work is sponsored by AMD’s High Performance Compute Fund project. This work is also partially sponsored by the Australian Research Council Discovery grant DP240102050, ARC LIEF grant LE240100131, and ARC Linkage grant LP230201022.
@@ D.1 Time Efficiency
Dataset Equation˜2 Equation˜4